# Trajectory learning for ensemble forecasts via the continuous ranked probability score: a Lorenz ’96 case study

### v1

@@ Abstract
This paper demonstrates the feasibility of trajectory learning for ensemble forecasts by employing the continuous ranked probability score (CRPS) as a loss function. Using the two-scale Lorenz ’96 system as a case study, we develop and train both additive and multiplicative stochastic parametrizations to generate ensemble predictions. Results indicate that CRPS-based trajectory learning produces parametrizations that are both accurate and sharp. The resulting parametrizations are straightforward to calibrate and outperform derivative-fitting-based parametrizations in short-term forecasts. This approach is particularly promising for data assimilation applications due to its accuracy over short lead times.
@@ 1 Introduction
Stochastic methods have been effectively utilized in numerical weather prediction for several decades [22, 5]. In recent years, data-driven stochastic models have been actively developed. The two-scale Lorenz ’96 system [32], a simplified representation of atmospheric dynamics with interactions between slow and fast scales, commonly serves as an initial test for novel calibration strategies [4]. Neglecting the fast variables is analogous to coarsening in fluid simulations, and hence introduces error and uncertainty in the evolution of the slow variables. Consequently, data from fast scales can inform stochastic parametrizations. For instance, [45] and [2] employed explicitly computed sub-grid scale data to tune parametrizations, fitting them as polynomials with a stochastic component for residuals, modeled as correlated or uncorrelated noise. Ensemble predictions with stochastic forcing were found to outperform perturbed parameter ensembles, with correlated additive noise yielding reliable forecasts. [10] modeled the sub-grid effects using Markov chains conditional on the resolved variables, achieving good agreement in the autocorrelation of the resolved variables and in their long-time distribution. Generative adversarial networks were used in [16] to construct maps between resolve variables and sub-grid scale forcings, producing accurate spatiotemporal correlations.
@@
In this work, we aim to produce a forecast with an CRPS and thereby tackle the model calibration challenge. While this challenge does not coincide with parameter estimation in stochastic differential equations (SDEs), they are related. Methods based upon maximum likelyhood estimation [39, 38], generalized method of moments [21], and the efficient method of moments [17] are among the approaches outlined in the review [37] that bear similarity to the CRPS learning approach adopted in this paper. However, parameter estimation in SDEs differs from producing a skillful forecast. The CRPS approach used in this study, should be viewed as both a data assimilation technique for ensemble forecasting and as a model calibration strategy [27].
Compensating for model error and unresolved dynamics remains a longstanding challenge in computational fluid dynamics, with the resulting models are generally referred to as ‘closure models’. Machine learning has proven to be an efficient approach for model calibration (see [41] and references therein). A robust method is trajectory fitting, where the discrepancy between predicted solution trajectories and reference trajectories is minimized [36]. However, many learned closure models are deterministic. Learning stochastic models is a logical next step necessary for probing prediction uncertainty via ensemble methods [6], and currently forms an active research area. A particular inspiration for the present study is the recent work by [30], who trained a stochastic model based on the almost fair CRPS to avoid problems with score degeneracy that can arise when all but one ensemble members have the same value as the reference. This approach is employed in the Artificial Intelligence Forecasting System (AIFS) developed at the ECMWF in test cases at two resolutions, where the trained stochastic model demonstrated improved variability compared to train deterministic models. The stochastic model was optimized for short-range forecasts but also showed good performance for longer range forecasts. Similarly, in [29] neural global circulation models for the atmosphere are trained using a stochastic loss based on the CRPS, giving strong practical evidence suggesting that minimising the CPRS to predict short-term weather is an effective way to tune parameterizations for climate.
@@ 2.1 Deterministic sub-grid scale measurement procedure
That is, the objective is to match the instantaneous coarse-grained dynamics over a given training data set. Here, $p$ determines how strongly outliers are penalized. Naturally, regularization terms can be included in the loss function without loss of generality, but they will not be further considered here.
An alternative approach is trajectory learning, where the model aims to minimize the distance between the low-fidelity dynamics and observed coarse-grained high-fidelity dynamics over specified integration intervals. We denote a reference trajectory by $q(t_{j}),\leavevmode\nobreak\ j=1,\ldots,N_{t}$, generated by the dynamics $\dot{q}=L(q)$, and a model trajectory by $z_{\theta}(t_{j}),\leavevmode\nobreak\ j=1,\ldots,N_{t}$, generated by the low-fidelity dynamics including the model as $\dot{z}=L_{c}(z)+M(z,\theta)$. Then, the model parameters $\theta$ are chosen such that the loss function
@@
is minimized over trajectories within a training data set. Here, $N_{t}$ denotes the length of each trajectory. Trajectory fitting can thus account for long-term dynamics (temporal non-locality) and error accumulation when the trajectory length is appropriately chosen.
A concise comparison between derivative fitting and trajectory learning is provided in [36, Table 1]. In brief, the difference between derivative fitting and trajectory learning decreases when the trajectory length is shortens and the time instances $t_{j}$ become sufficiently close. However, when data are sparse or noisy, approximations of the instantaneous dynamics based on available measurements become less accurate and derivative fitting loses reliability. Simultaneously, derivative fitting requires time derivatives of training data, which may not always be readily available. Conversely, trajectory fitting requires either a discretized adjoint equation or a differentiable numerical solver. Furthermore, if the trajectory length is too long or the time instances are spaced too far apart, chaotic behavior and perturbation sensitivity of the dynamical system may decrease the effectiveness of trajectory learning.
@@ 2.2 Probabilistic sub-grid scale measurement procedure
The distribution $\pi_{\mathrm{CG}}^{t_{1}}$ of coarse-grained field will generally not be known. Monte Carlo approximations can be computed using the high-fidelity flow map $\psi_{[t_{0},t_{1}]}$, yet this map can be unknown or may be prohibitively expensive to compute. Instead, a single observed trajectory can be treated as a time-varying delta distribution to facilitate probabilistic trajectory learning. To this end, we consider an ensemble forecast $\{z_{\theta}^{i}(t_{j})\}_{i=1,\ldots,M},\leavevmode\nobreak\ j=1,\ldots,N_{t}$, generated by the flow map (6). Here, the superscript $i$ denotes the $i^{\mathrm{th}}$ ensemble member, $M$ is the ensemble size, and $N_{t}$ is the trajectory length. As before, we let $q(t_{j}),\leavevmode\nobreak\ j=1,\ldots,N_{t},$ be the reference trajectory generated by the high-fidelity dynamics. The continuous ranked probability score (CRPS) then serves as a probabilistic extension to the mean absolute error, enabling trajectory learning for ensemble predictions. The objective becomes to choose model parameters $\theta$ such that the loss function
$\begin{split}\mathrm{Loss}(\theta,q)&=\sum_{j=1}^{N_{t}}\mathrm{CRPS}(z_{\theta},C(q),t_{j})\\ \mathrm{CRPS}(z_{\theta},q,t_{j})&=\frac{1}{M}\sum_{i=1}^{M}\left|z_{\theta}^{i}(t_{j})-C(q(t_{j}))\right|+\frac{1}{2M^{2}}\sum_{j=1}^{M}\sum_{k=1}^{M}\left|z_{\theta}^{i}(t_{j})-z_{\theta}^{k}(t_{j})\right|\end{split}$ (7)
is minimized over a set of training trajectories. A small CRPS implies both an accurate ensemble mean absolute error and a sharp ensemble forecast (small spread). Furthermore, the CRPS is suitable for ensemble forecasts of nonlinear systems, as no assumptions are made on the distribution of the samples. As a proper scoring rule, the CRPS is minimized only when the forecast distribution matches the observed distribution. This ensures that training with the aforementioned loss function favors both ensemble accuracy and sharpness. Furthermore, the absence of squared errors in the loss function makes the training less sensitive to large outliers, thereby increasing training robustness in chaotic systems. For additional details regarding the CRPS, we refer to [3, 18, 19, 23, 35]. An illustration of CRPS-based trajectory learning of ensemble forecasts is provided in Fig. 1.
@@ 3 Governing equations and models
In this section, we detail the governing equations and stochastic parametrizations. We differentiate between ‘true’ reference dynamics, given by the deterministic two-scale Lorenz ’96 (L96) model, integrated at high temporal resolution. A computationally cheaper ‘coarsened’ forward model comprises a single-scale L96 model, integrated at low temporal resolution. Thus, the latter exhibits deficiencies due to reduced temporal resolution and the omission of fast-scale dynamics. These deficiencies are parametrized by means of stochastic models, using synthetic data generated from the high-fidelity reference dynamics.
@@ 3.1 Governing equations
The two parameter sets considered in this study are identical to those presented by [2] and are summarized in Table 1. Time is expressed in terms of model time units (MTU), where one MTU corresponds to approximately 5 atmospheric days.
@@
The high-fidelity solution is initialized by assigning to all variables an independent sample from a standard normal distribution, after which the solution is integrated for 1500 MTU to remove any transitory behavior. The obtained solution at this time serves as the starting point for the data measurement procedure. Subsequently, the truth is simulated for an additional 500 MTU to generate training and verification data.
@@ 3.2.1 Local model
The stochastic components of the parametrization follow from the residuals of the polynomial fit. Specifically, two additive stochastic processes are considered, based on the sample variance $S^{2}$ of the residuals and their autocorrelation $\phi$. The first process consists of uncorrelated, normally distributed noise, where the increments over a single time step are drawn from $\mathcal{N}(0,S^{2}\Delta t)$. The second process is an Ornstein–Uhlenbeck (OU) process that takes into account measured autocorrelation of the residuals. It is discretized as a zero-mean first-order autoregressive (AR(1)) process, which evolves as
$r(t+\Delta t)=\phi\hskip 1.42262ptr(t)+\sqrt{\Delta tS^{2}(1-\phi^{2})}\leavevmode\nobreak\ \varepsilon(t).$ (13)
Here, $\varepsilon(t)\sim\mathcal{N}(0,1)$ and the scaling term ensures that the resulting samples have the same variance as the uncorrelated model. The value of $\phi$ is set to the autocorrelation of the residuals after $\Delta t=0.005$ MTU. All parameter values are summarized in Table 2.
@@ 3.2.2 Global model
Figure 2: POD modes for $c=4$ (left) and $c=10$ (right) that serve as the basis for the global stochastic parametrization. The pairs of POD modes are shown with a vertical offset for visibility, all modes have zero mean. The mean profile $\xi_{0}$ is not offset.
@@ 3.3 CRPS-based forecast model
and setting $c_{i}(X)=1$. Here $\mathbf{r}\in\mathbb{R}^{K}$ denotes the value of the stochastic process, $\boldsymbol{\mu}\in\mathbb{R}^{K}$ is the mean, $\mathbf{A}\in\mathbb{R}^{K\times K}$ defines the coupling between the different components, $\mathbf{B}\in\mathbb{R}^{K\times q}$ determines the noise scaling, and $\boldsymbol{\varepsilon}\in\mathbb{R}^{q}$ are $q$ independent Brownian motions, that is, $\varepsilon_{i}(t)\sim N(0,1)$. In what follows, we choose $q=K$. The coupled OU process offers greater flexibility compared to the previously used independent scalar OU processes. Namely, interaction between the modes is encoded in $\mathbf{A}$, and incorporating the mean $\boldsymbol{\mu}$ removes the need to include the mean profile $\xi_{0}$ into the model.
@@ 3.3.1 Model training
The model is trained during 400 epochs, after which the loss per batch was found to have stabilized. Each epoch consists of 125 batches. The Adam optimizer [28] is used with a fixed learning rate set at $10^{-2}$ for training the additive noise model and $10^{-4}$ for the multiplicative noise model. A series of trajectory lengths is used, ranging from $N_{t}=2$ (0.01 MTU) to $N_{t}=200$ (1 MTU). A full description of the batch size, trajectory length, and final measured loss per batch is provided in Appendix A.
The loss per batch during training is presented in Fig. 3 for different values of $N_{t}$. We observe that a short trajectory length for the additive noise model yields only a marginal decrease in the loss. Since each training trajectory starts from a single perfect initial condition, the error and spread after two time steps are small. Hence, for such short time intervals, the initial loss shows is barely improved upon during training. This effect diminishes for longer trajectories, and a decrease of the loss during training is realized until the trajectories become too long. Unstable training behavior is observed for $N_{t}=128$ and $N_{t}=200$, which is attributed to the chaotic dynamics of the system. Specifically, long model trajectories will at some point diverge from the reference trajectories. Minor changes in model parameters will induce significant changes in trajectory due to high system sensitivity, and too much effort is put in correcting long lead-time dynamics. This can lead to unstable gradients and inefficient optimization. This issue could be mitigated by weighting the loss function such that long lead-time predictions contribute less to the loss [36]; however, this is not further investigated here.
The evolution of the loss for the multiplicative noise model differs from that of the additive noise model. We observe that a short trajectory length ($N_{t}=2$) leads to a significant, rapid decrease of the loss, owing to the increased model parameters when compared to the additive noise model. Furthermore, training becomes less effective even at modest trajectory lengths, suggesting higher sensitivity to the model parameters as trajectory length increases.
Figure 3: Evolution of the loss per batch over the training duration, for $c=4$ (left column) and $c=10$ (right column), for the additive noise model (top row) and the multiplicative noise model (bottom row). The displayed values are normalized by the initial value for visual comparison, and a moving average is applied for clarity.

### v2

@@ Abstract
This paper demonstrates the feasibility of trajectory learning for ensemble forecasts by employing the continuous ranked probability score (CRPS) as a loss function. Using the two-scale Lorenz ’96 system as a case study, we develop and train both additive and multiplicative stochastic parametrizations to generate ensemble predictions. Results indicate that CRPS-based trajectory learning produces parametrizations that are both accurate and sharp, while the self-spread term in the loss function regularises training. The resulting parametrizations are straightforward to calibrate and outperform derivative-fitting-based parametrizations in short-term forecasts. This approach is particularly promising for data assimilation applications due to its accuracy over short lead times.
@@ 1 Introduction
Stochastic methods have been effectively utilised in numerical weather prediction for several decades [22, 5]. In recent years, data-driven stochastic models have been actively developed. The two-scale Lorenz ’96 system [32], a simplified representation of atmospheric dynamics with interactions between slow and fast scales, commonly serves as an initial test for novel calibration strategies [4]. Neglecting the fast variables is analogous to coarsening in fluid simulations, and hence introduces error and uncertainty in the evolution of the slow variables. Consequently, data from fast scales can inform stochastic parametrizations. For instance, [45] and [2] employed explicitly computed sub-grid scale data to tune parametrizations, fitting them as polynomials with a stochastic component for residuals, modeled as correlated or uncorrelated noise. Ensemble predictions with stochastic forcing were found to outperform perturbed parameter ensembles, with correlated additive noise yielding reliable forecasts. [10] modeled the sub-grid effects using Markov chains conditional on the resolved variables, achieving good agreement in the autocorrelation of the resolved variables and in their long-time distribution. Generative adversarial networks were used in [16] to construct maps between resolved variables and sub-grid scale forcings, producing accurate spatiotemporal correlations.
@@
In this work, we aim to produce a forecast with an CRPS and thereby tackle the model calibration challenge. While this challenge does not coincide with parameter estimation in stochastic differential equations (SDEs), they are related. Methods based upon maximum likelyhood estimation [39, 38], generalised method of moments [21], and the efficient method of moments [17] are among the approaches outlined in the review [37] that bear similarity to the CRPS learning approach adopted in this paper. However, parameter estimation in SDEs differs from producing a skillful forecast. The CRPS approach used in this study, can be viewed as both a data assimilation technique for ensemble forecasting and as a model calibration strategy [27].
Compensating for model error and unresolved dynamics remains a longstanding challenge in computational fluid dynamics, with the resulting models generally referred to as ‘closure models’. Machine learning has proven to be an efficient approach for model calibration (see [41] and references therein). A robust method is trajectory fitting, where the discrepancy between predicted solution trajectories and reference trajectories is minimised [36]. However, many learned closure models are deterministic. Learning stochastic models is a logical next step necessary for probing prediction uncertainty via ensemble methods [6], and currently forms an active research area. A particular inspiration for the present study is the recent work by [30], who trained a stochastic model based on the almost fair CRPS to avoid problems with score degeneracy that can arise when all but one ensemble member has the same value as the reference. This approach is employed in the Artificial Intelligence Forecasting System (AIFS) developed at the ECMWF in test cases at two resolutions, where the trained stochastic model demonstrated improved variability compared to train deterministic models. The stochastic model was optimised for short-range forecasts but also showed good performance for longer range forecasts. Similarly, in [29] neural global circulation models for the atmosphere are trained using a stochastic loss based on the CRPS, giving strong practical evidence suggesting that minimising the CPRS to predict short-term weather is an effective way to tune parameterizations for climate.
@@ 2.1 Deterministic sub-grid scale measurement procedure
That is, the objective is to match the instantaneous coarse-grained dynamics over a given training data set. Here, $p$ determines how strongly outliers are penalised. Naturally, regularization terms can be included in the loss function without loss of generality, but they will not be further considered here.
An alternative approach is trajectory learning, where the model aims to minimize the distance between the low-fidelity dynamics and observed coarse-grained high-fidelity dynamics over specified integration intervals. We denote a reference trajectory by $q(t_{j}),\penalty 10000\ j=1,\ldots,N_{t}$, generated by the dynamics $\dot{q}=L(q)$, and a model trajectory by $z_{\theta}(t_{j}),\penalty 10000\ j=1,\ldots,N_{t}$, generated by the low-fidelity dynamics including the model as $\dot{z}=L_{c}(z)+M(z,\theta)$. Then, the model parameters $\theta$ are chosen such that the loss function
@@
is minimised over trajectories within a training data set. Here, $N_{t}$ denotes the length of each trajectory. Trajectory fitting can thus account for long-term dynamics (temporal non-locality) and error accumulation when the trajectory length is appropriately chosen.
A concise comparison between derivative fitting and trajectory learning is provided in [36, Table 1]. In brief, the difference between derivative fitting and trajectory learning decreases when the trajectory length is shortens and the time instances $t_{j}$ become sufficiently close. However, when data are sparse or noisy, approximations of the instantaneous dynamics based on available measurements become less accurate and derivative fitting loses reliability. Simultaneously, derivative fitting requires time derivatives of training data, which may not always be readily available. Conversely, trajectory fitting requires either a discretised adjoint equation or a differentiable numerical solver. Furthermore, if the trajectory length is too long or the time instances are spaced too far apart, chaotic behavior and perturbation sensitivity of the dynamical system may decrease the effectiveness of trajectory learning.
@@ 2.2 Probabilistic sub-grid scale measurement procedure
The distribution $\pi_{\mathrm{CG}}^{t_{1}}$ of coarse-grained field will generally not be known. Monte Carlo approximations can be computed using the high-fidelity flow map $\psi_{[t_{0},t_{1}]}$, yet this map can be unknown or may be prohibitively expensive to compute. Instead, a single observed trajectory can be treated as a time-varying delta distribution to facilitate probabilistic trajectory learning. To this end, we consider an ensemble forecast $\{z_{\theta}^{i}(t_{j})\}_{i=1,\ldots,M},\penalty 10000\ j=1,\ldots,N_{t}$, generated by the flow map (6). Here, the superscript $i$ denotes the $i^{\mathrm{th}}$ ensemble member, $M$ is the ensemble size, and $N_{t}$ is the trajectory length. As before, we let $q(t_{j}),\penalty 10000\ j=1,\ldots,N_{t},$ be the reference trajectory generated by the high-fidelity dynamics. The continuous ranked probability score (CRPS) then serves as a probabilistic extension to the mean absolute error, enabling trajectory learning for ensemble predictions. The objective becomes to choose model parameters $\theta$ such that the loss function
$\begin{split}\mathrm{Loss}(\theta,q)&=\sum_{j=1}^{N_{t}}\mathrm{CRPS}(z_{\theta},C(q),t_{j}),\\ \mathrm{CRPS}(z_{\theta},q,t_{j})&=\frac{1}{M}\sum_{i=1}^{M}\left|z_{\theta}^{i}(t_{j})-C(q(t_{j}))\right|-\frac{1}{2M^{2}}\sum_{j=1}^{M}\sum_{k=1}^{M}\left|z_{\theta}^{i}(t_{j})-z_{\theta}^{k}(t_{j})\right|,\end{split}$ (7)
is minimised over a set of training trajectories.
The first term in the CRPS represents the ensemble mean absolute error, while the second term quantifies the ensemble spread and is subtracted from the mean error. On its own, a larger spread would reduce the CRPS. However, in practice, an increase in spread usually leads to a larger mean error, which causes the CRPS to increase. Hence, a small CRPS implies both an accurate ensemble mean absolute error and a sharp ensemble forecast (small spread). Furthermore, the CRPS is suitable for ensemble forecasts of nonlinear systems, as no assumptions are made on the distribution of the samples. As a proper scoring rule, the CRPS is minimised only when the forecast distribution matches the observed distribution. This ensures that training with the aforementioned loss function favors both ensemble accuracy and sharpness. Moreover, the absence of squared errors in the loss function makes the training less sensitive to large outliers, thereby increasing training robustness in chaotic systems. For additional details regarding the CRPS, we refer to [3, 18, 19, 23, 35]. An illustration of CRPS-based trajectory learning of ensemble forecasts is provided in Fig. 1.
@@ 3 Governing equations and models
In this section, we detail the governing equations and stochastic parametrizations. We differentiate between ‘true’ reference dynamics, given by the deterministic two-scale Lorenz ’96 (L96) model, integrated at high temporal resolution. A computationally cheaper ‘coarsened’ forward model comprises a single-scale L96 model, integrated at low temporal resolution. Thus, the latter exhibits deficiencies due to reduced temporal resolution and the omission of fast-scale dynamics. These deficiencies are parametrised by means of stochastic models, using synthetic data generated from the high-fidelity reference dynamics.
@@ 3.1 Governing equations
The two parameter sets considered in this study are identical to those presented by [2] and are summarised in Table 1. Time is expressed in terms of model time units (MTU), where one MTU corresponds to approximately 5 atmospheric days.
@@
The high-fidelity solution is initialised by assigning to all variables an independent sample from a standard normal distribution, after which the solution is integrated for 1500 MTU to remove any transitory behavior. The obtained solution at this time serves as the starting point for the data measurement procedure. Subsequently, the truth is simulated for an additional 500 MTU to generate training and verification data.
@@ 3.2.1 Local model
The stochastic components of the parametrization follow from the residuals of the polynomial fit. Specifically, two additive stochastic processes are considered, based on the sample variance $S^{2}$ of the residuals and their autocorrelation $\phi$. The first process consists of uncorrelated, normally distributed noise, where the increments over a single time step are drawn from $\mathcal{N}(0,S^{2}\Delta t)$. The second process is an Ornstein–Uhlenbeck (OU) process that takes into account measured autocorrelation of the residuals. It is discretised as a zero-mean first-order autoregressive (AR(1)) process, which evolves as
$r(t+\Delta t)=\phi\hskip 1.42262ptr(t)+\sqrt{\Delta tS^{2}(1-\phi^{2})}\penalty 10000\ \varepsilon(t).$ (13)
Here, $\varepsilon(t)\sim\mathcal{N}(0,1)$ and the scaling term ensures that the resulting samples have the same variance as the uncorrelated model. The value of $\phi$ is set to the autocorrelation of the residuals after $\Delta t=0.005$ MTU. All parameter values are summarised in Table 2.
@@ 3.2.2 Global model
Figure 2: POD modes for $c=4$ (left) and $c=10$ (right) that serve as the basis for the global stochastic parametrization. The pairs of POD modes are shown with a vertical offset for visibility, all modes have zero mean. The mean profile $\xi_{0}$ does not have zero mean and is not offset.
@@ 3.3 CRPS-based forecast model
and setting $c_{i}(X)=1$. Here $\mathbf{r}\in\mathbb{R}^{K}$ denotes the value of the stochastic process, $\boldsymbol{\mu}\in\mathbb{R}^{K}$ is the mean, $\mathbf{A}\in\mathbb{R}^{K\times K}$ defines the coupling between the different components, $\mathbf{B}\in\mathbb{R}^{K\times q}$ determines the noise scaling, and $\boldsymbol{\varepsilon}\in\mathbb{R}^{q}$ are $q$ independent Brownian motions, that is, $\varepsilon_{i}(t)\sim N(0,1)$. In what follows, we choose $q=K$. The coupled OU process offers greater flexibility compared to the previously used independent scalar OU processes. Namely, interaction between the modes is encoded in $\mathbf{A}$, and incorporating the mean $\boldsymbol{\mu}$ removes the need to include the mean profile $\xi_{0}$ in the model.
@@ 3.3.1 Model training
The model is trained during 400 epochs, after which the loss per batch was found to have stabilised. Each epoch consists of 125 batches. The Adam optimizer [28] is used with a fixed learning rate set at $10^{-2}$ for training the additive noise model and $10^{-4}$ for the multiplicative noise model. A series of trajectory lengths is used, ranging from $N_{t}=2$ (0.01 MTU) to $N_{t}=200$ (1 MTU). A full description of the batch size, trajectory length, and final measured loss per batch is provided in Appendix A.
The loss per batch during training is presented in Fig. 3 for different values of $N_{t}$. We observe that a short trajectory length for the additive noise model yields only a marginal decrease in the loss. Since each training trajectory starts from a single perfect initial condition, the error and spread after two time steps are small. Hence, for such short time intervals, the initial loss shows is barely improved upon during training. This effect diminishes for longer trajectories, and a decrease of the loss during training is realised until the trajectories become too long. Unstable training behavior is observed for $N_{t}=128$ and $N_{t}=200$, which is attributed to the chaotic dynamics of the system. Specifically, long model trajectories will at some point diverge from the reference trajectories. Minor changes in model parameters will induce significant changes in trajectory due to high system sensitivity, and too much effort is put in correcting long lead-time dynamics. This can lead to unstable gradients and inefficient optimization. This issue could be mitigated by weighting the loss function such that long lead-time predictions contribute less to the loss [36]; however, this is not further investigated here.
The evolution of the loss for the multiplicative noise model differs from that of the additive noise model. We observe that a short trajectory length ($N_{t}=2$) leads to a significant, rapid decrease of the loss, owing to the increased model parameters when compared to the additive noise model. Furthermore, training becomes less effective even at modest trajectory lengths, suggesting higher sensitivity to the model parameters as the trajectory length increases.
Figure 3: Evolution of the loss per batch over the training duration, for $c=4$ (left column) and $c=10$ (right column), for the additive noise model (top row) and the multiplicative noise model (bottom row). The displayed values are normalised by the initial value for visual comparison, and a moving average is applied for clarity.